# Solving Inverse Problems with Model Mismatch using Untrained Neural Networks within Model-based Architectures

### v1

@@ Abstract
Model-based deep learning methods such as loop unrolling (LU) and deep equilibrium model (DEQ) extensions offer outstanding performance in solving inverse problems (IP). These methods unroll the optimization iterations into a sequence of neural networks that in effect learn a regularization function from data. While these architectures are currently state-of-the-art in numerous applications, their success heavily relies on the accuracy of the forward model. This assumption can be limiting in many physical applications due to model simplifications or uncertainties in the apparatus. To address forward model mismatch, we introduce an untrained forward model residual block within the model-based architecture to match the data consistency in the measurement domain for each instance. We propose two variants in well-known model-based architectures (LU and DEQ) and prove convergence under mild conditions. The experiments show significant quality improvement in removing artifacts and preserving details across three distinct applications, encompassing both linear and nonlinear inverse problems. Moreover, we highlight reconstruction effectiveness in intermediate steps and showcase robustness to random initialization of the residual block and a higher number of iterations during evaluation.
@@ 1 Introduction
where $\gamma\geq 0$ is an appropriately-chosen parameter. The regularizer $r$ is usually predetermined based on some known or desired structure, e.g., $\ell_{1}$-, $\ell_{2}$-, or total variation (TV) norm to promote sparsity, smoothness, or edges in image reconstructions, respectively. Solving (2) requires careful consideration of the underlying physics or the forward model $\mathcal{A}$ to obtain a stable and accurate reconstruction. However, knowing $\mathcal{A}$ can be challenging in practice. The reasons for this include inaccurate measurements and challenging calibration, highly nonlinear and/or computationally expensive models replaced by simplified versions, as well as access to only approximations of certain features. Usually, some knowledge of the true model, designated $\mathcal{A}_{0}$ in this work, is available. This occurs in many applications, including blind deconvolution/deblurring problems, recovering seismic layer models using the simplified acoustic wave equation as the forward model (Mousa & Al-Shuhail, 2011), determining fault locations in media with unknown structures (Mahmoud & Khalid, 2013), etc.
@@
Some works passively address errors in the $\bm{y}$-space (equivalent to model mismatch under some conditions) by enhancing the robustness of the neural network. For example, Krainovic et al. (2023) suggests that introducing jittering in the $\bm{y}$-space can achieve robustness in worst-case $\ell_{2}$ perturbations. In Hu et al. (2023), a deep equilibrium solver is trained using various incorrect $\mathcal{A}$’s, demonstrating greater robustness to variations in $\mathcal{A}$ compared to the Plug-and-Play method. In contrast, our proposed approach involves an active strategy for mitigating model mismatch. The experimental results in Section 5 illustrate the effectiveness of our method compared to the same LU network that merely learns a robust mapping passively.
Model mismatches in linear inverse problems can be more directly resolved by alternatively updating the model parameters and the underlying signal. For instance, Fergus et al. (2006); Cai et al. (2009); Levin et al. (2009); Cho & Lee (2009) reconstructed the latent images with updates in the forward model solved directly from least squares. However, these methods are typically limited to linear cases with predefined regularizers. To improve performance in this setting, Nan & Ji (2020); Gilton et al. (2021b) learn a regularizer from data, resulting in improved performance on linear inverse problems. Our work further introduces an untrained residual network that can learn the forward model more flexibly to handle nonlinear problems and hence has broader scientific applications. Existing LU approaches learn a black-box solver that treats the provided $\mathcal{A}_{0}$ as if it is the true forward model, while the proposed method updates the forward model and the reconstruction jointly, providing both a solution to the IP and an estimate of the true forward model.
Our contributions in this paper can be summarized as follows:
@@
  • General model-based architecture in handling model-mismatch: We present a general model-based architecture for solving IPs when an approximation of the forward model is available. Unlike classical model-based architectures which require precise knowledge of $\mathcal{A}$, we propose two variants of the model-based algorithms that can iteratively update the forward model along with the reconstruction.
@@
  • Introducing untrained neural network for model-mismatch: We present a novel approach by incorporating an untrained neural network that can match the data consistency in the measurement domain within a model-based architecture to accommodate model-mismatch in both linear and nonlinear IPs. We demonstrate that a random initialization of this network during evaluation maintains the same level of reconstruction.
@@
  • Improved performance in linear and nonlinear IP tasks: In contrast to model-based architectures passively trained to be robust to model mismatch, our methods exhibit substantial improvement in image blind deblurring (linear), seismic blind deconvolution (linear), and landscape defogging (nonlinear) tasks.
@@ 5 Experiments and Discussion
where $\bm{T}\in[0,1]^{3\times m\times n}$ is an unknown transmission map as a function of the fog-free image $\bm{x}$, $m$ and $n$ denote the width and height of the image. $\bm{T}(\bm{x})$ is then applied pointwise ($\odot$) to $\bm{x}$. $\bm{L}(\bm{x})$ is the unknown atmospheric light profile (Tufail et al., 2019). The goal is to estimate the model mismatch and recover the clean image. Since $\bm{T}$ and $\bm{L}$ are unknown functions with respect to $\bm{x}$, we use two separate networks to learn transmission mapping and the light profile respectively. This example shows the flexibility of our proposed methods. Instead of having one single network $f_{\theta}$ that learns the residual, if we know the explicit formulation of the true model, we can customize the residual network in a more structured way. To evaluate the proposed approachs, the CityScapes - Depth and Segmentation dataset given in Cordts et al. (2016) is used. This dataset provides depth maps, as well as foggy and clean views of urban street scenes.
5.1 Training Details
@@ 5.2 Reconstruction Results
Table 1 compares the average testing peak signal-to-noise ratio (PSNR) in dB and Structural Similarity Index Measure (SSIM) of the reconstructions. The neural network learns a direct inverse regardless of $\mathcal{A}_{0,i}$, which is the worst in reconstruction. It is notable that, even with some degree of inaccuracy in $\mathcal{A}$, the robust LU still outperforms the direct inverse mapping. This is because end-to-end training enables LU to fix the gradient-step errors resulting from the model mismatch to some extent, thereby maintaining reasonably higher-quality results. Our results align with the finding in Krainovic et al. (2023) that the robust LU trains an inverse mapping that is less sensitive to perturbations in the $\bm{y}$-space, by seeing inaccuracies in the forward model at train time. However, in the next section, we will discuss the intermediate reconstruction ineffectiveness of robust LU compared to the proposed methods.
@@
The proposed methods make an improvement by actively adjusting for the model mismatch for each data pair. Instead of having the neural network learn both the proximal step and the model correction as in a black box, the proposed methods separate the tasks and each part of the network is now learning a clearer and easier task, resulting in a substantial performance gain over both the robust LU and the direct inverse mapping in all tasks. Furthermore, $\mathcal{A}$-adaptive DEQ allows for a higher number of iterations than $\mathcal{A}$-adaptive LU, resulting in superior performance in both PSNR and SSIM.
Table 1: Average testing PSNR and SSIM for direct inverse mapping, robust LU, $\mathcal{A}$-adaptive LU, and $\mathcal{A}$-adaptive DEQ. The best performances for each task are in bold and the second best performances are underlined.
@@
Direct Inverse PSNR 24.002 22.816 21.802
SSIM 0.7689 0.7421 0.8738
Robust LU PSNR 34.380 24.785 29.645
SSIM 0.9429 0.8407 0.9563
$\mathcal{A}$-adaptive LU PSNR 36.621 27.271 31.112
(proposed meth. I) SSIM 0.9731 0.8960 0.9661
$\mathcal{A}$-adaptive DEQ PSNR 37.809 28.236 32.149
(proposed meth. II) SSIM 0.9774 0.9082 0.9723
@@ 5.4 Intermediate Reconstructions of $\mathcal{A}$-adaptive DEQ
We show the intermediate reconstructions of the DEQ variant of the proposed method in Fig. 7. The proposed network is trained with a maximum of 30 iterations or when a fixed point solution is reached, whereas when evaluating $\mathcal{A}$-adaptive DEQ with more iterations (i.e., 100), the reconstruction error remains low. This result aligns with the finding in Gilton et al. (2021a).
@@ 5.5 Robustness of $\mathcal{A}$-adaptive LU with More Iterations
In this section, we empirically assess the stability of our proposed $\mathcal{A}$-adaptive LU variant by evaluating it with a greater number of unrolling iterations than those employed during training. Across all three tasks, a noticeable Mean Squared Error (MSE) gap emerges between the two curves. The solid blue curve, representing robust LU, exhibits a drop in MSE at iteration 5, which corresponds to the target output during training. However, as the number of iterations exceeds the max iterations of training (in this set of experiments $k>5$), the MSE rises significantly. Notice that although the MSE at iteration 10 is relatively low for the landscape defogging task, the reconstruction is not stable and less predictable. In contrast, the dashed red curve ($\mathcal{A}$-adaptive LU) oscillates around the lowest MSE. This observation highlights the enhanced robustness of the proposed $\mathcal{A}$-adaptive LU when subjected to a higher number of iterations during evaluation.
@@ 6 Conclusion
This work adaptively addresses the errors in the forward model by optimizing the auxiliary variable and the untrained mismatch neural network within a model-based architecture. We show the flexibility of the proposed forward model matching module using two well-known architectures. However, the accuracy-runtime tradeoff still exists. Although the proposed $\mathcal{A}$-adaptive DEQ variant achieves the best performance for all tasks, when solving large-scale inverse problems, careful consideration of runtime is necessary.

### v2

@@ Abstract
Model-based deep learning methods such as loop unrolling (LU) and deep equilibrium model (DEQ) extensions offer outstanding performance in solving inverse problems (IP). These methods unroll the optimization iterations into a sequence of neural networks that in effect learn a regularization function from data. While these architectures are currently state-of-the-art in numerous applications, their success heavily relies on the accuracy of the forward model. This assumption can be limiting in many physical applications due to model simplifications or uncertainties in the apparatus. To address forward model mismatch, we introduce an untrained forward model residual block within the model-based architecture to match the data consistency in the measurement domain for each instance. We propose two variants in well-known model-based architectures (LU and DEQ) and prove convergence under mild conditions. Our approach offers a unified solution that is less parameter-sensitive, requires no additional data, and enables simultaneous fitting of the forward model and reconstruction in a single pass, benefiting both linear and nonlinear inverse problems. The experiments show significant quality improvement in removing artifacts and preserving details across three distinct applications, encompassing both linear and nonlinear inverse problems. Moreover, we highlight reconstruction effectiveness in intermediate steps and showcase robustness to random initialization of the residual block and a higher number of iterations during evaluation. Code is available at https://github.com/InvProbs/A-adaptive-model-based-methods.
@@ 1 Introduction
where $\gamma\geq 0$ is an appropriately-chosen parameter. The regularizer $r$ is usually predetermined based on some known or desired structure, e.g., $\ell_{1}$-, $\ell_{2}$-, or total variation (TV) norm to promote sparsity, smoothness, or edges in image reconstructions, respectively. Solving (2) requires careful consideration of the underlying physics or the forward model $\mathcal{A}$ to obtain a stable and accurate reconstruction. However, knowing $\mathcal{A}$ can be challenging in practice. The reasons for this include inaccurate measurements and challenging calibration, highly nonlinear and/or computationally expensive models replaced by simplified versions, as well as access to only approximations of certain features. Usually, some knowledge of the true model, designated $\mathcal{A}_{0}$ in this work, is available. This occurs in many applications, including blind deconvolution/deblurring problems, recovering seismic layer models using the simplified acoustic wave equation as the forward model (Mousa & Al-Shuhail, 2011), determining fault locations in media with unknown structures (Mahmoud & Khalid, 2013), computational tomography (CT) when the dynamic behavior of a human object is observed during the measurement period Blanke et al. (2020), sparse signal recovery when the sensing matrix is not known perfectly, etc.
@@
Some works passively address errors in the $\bm{y}$-space by enhancing the robustness of the neural network. For example, Krainovic et al. (2023) suggests that introducing jittering in the $\bm{y}$-space can achieve robustness in worst-case $\ell_{2}$ perturbations. In Hu et al. (2023), a deep equilibrium solver is trained using various incorrect $\mathcal{A}$’s, demonstrating greater robustness to variations in $\mathcal{A}$ compared to the Plug-and-Play method. However, it is noteworthy that such robust networks often experience a tradeoff between accuracy and stability, as discussed in Krainovic et al. (2023); Gottschling et al. (2023). In contrast, our proposed approach involves an active strategy for mitigating model mismatch. The experimental results in Section 6 illustrate the effectiveness of our method compared to the same LU network that merely learns a robust mapping passively.
Various active forward model matching algorithms are tailored to different tasks. For linear IPs, model mismatches can be more easily resolved by alternatively updating the model parameters and the underlying signal. For instance, Fergus et al. (2006); Cai et al. (2009); Levin et al. (2009); Cho & Lee (2009) reconstructed the latent images with updates in the forward model solved directly from least squares. However, these methods are often constrained to linear scenarios and rely on task-specific recovery techniques, which are typically highly sensitive to hyperparameters Cho & Lee (2009). Later works Nan & Ji (2020); Gilton et al. (2021b) learn a regularizer from data, resulting in improved performance on linear inverse problems. The structure of the initial forward model in a nonlinear inverse problem often differs from the true model. For instance, linear models may be used to estimate nonlinear functions. This motivates the use of neural networks to learn the true forward model (Candiani et al., 2021; Koponen et al., 2021; Lunz et al., 2021). However, traditional neural network training approaches require separate stages for learning the forward model and signal reconstruction. This requires a substantial amount of training data to ensure accurate mapping (Arjas et al., 2023) from any ${\bm{x}}$ (which represents ground truth and intermediate reconstructions) to its corresponding ${\bm{y}}$ in iterative reconstructions. With limited training data, this method falls out of the discussion.
In contrast, we introduce an untrained residual network capable of handling forward model mismatch for general IPs. The proposed method is a unified framework that applies to both linear and nonlinear IPs, offering enhanced stability during training, requiring no additional data, and allowing fitting the forward model and reconstruction simultaneously in a single pass. We further list our contributions as follows:
@@
  • General model-based architecture in handling model-mismatch: We present a general model-based architecture for solving IPs when only an approximation of the forward model is available. Unlike classical model-based architectures which require precise knowledge of $\mathcal{A}$, we propose two variants of the model-based algorithms that can iteratively update the forward model along with the reconstruction.
@@
  • Introducing untrained neural network for model-mismatch: We introduce a novel approach by integrating an untrained neural network into a model-based architecture to handle model-mismatch in both linear and nonlinear inverse problems. By updating its parameters alongside the reconstruction process, we circumvent the need for additional data and pre-training the forward model in solving nonlinear inverse problems. Indeed, we show that random initialization of this network during evaluation yields the same level of reconstruction.
@@
  • Improved performance in linear and nonlinear IP tasks: In contrast to the passive robust training approaches within general IP solvers using model-based architectures, our methods demonstrate substantial improvement in image blind deblurring (linear), seismic blind deconvolution (linear), and landscape defogging (nonlinear) tasks. It’s worth noting that while our application showcases improvements in these specific areas, its scope extends beyond to fields like medical imaging.
@@
2 Related Works
2.1 Review of Model-based Architectures
Model-based architectures in solving inverse problems are motivated by classical optimization. When the forward operator $\mathcal{A}$ is precisely known, one can unroll the objective function in (1) into numbers of iterative steps, and learn the regularization update from data. One subclass of model-based IP solvers is called loop unrolling (LU) or deep unfolding, which has a fixed number of iterations. LU is first introduced to solve sparse coding Gregor & LeCun (2010); Chen et al. (2018), and later on extended to many image restoration problems Monga et al. (2020); Gilton et al. (2021a); Ongie et al. (2020); Zhao et al. (2023; 2024). Variations of LU emerge according to different optimization algorithms, such as projected gradient descent Lohit et al. (2019), proximal gradient descent Hosseini et al. (2020), alternating direction method of multipliers (ADMM) Yang et al. (2020), and momentum gradient descent Chun et al. (2023). More works can be found in Monga et al. (2020).
The concept of unrolling optimization iterations is not restricted to a fixed number but is often limited by memory constraints during backpropagation. To address this, Gilton et al. (2021a) proposes to train the iterative updates of the reconstructions until convergence using a fixed point solver. Similarly, Cai et al. (2021) extends a loop unrolling to an infinite number of iterations for robust principal component analysis, achieved by layer-wise training the initial iterations and updating the additional penalty parameters for subsequent iterations. In our work, we demonstrate the proposed forward model matching technique using DEQ, known for its infinite-layer extension with consistent architecture across all iterations.
2.2 Methods for Addressing Model Mismatch
A linear inverse problem with forward model mismatch can be interpreted as a parameter tuning within a predetermined structure. In other words, given a matrix $\bm{A}$ of dimension $m\times n$, the entries contain errors. This happens in many imaging tasks, such as blind deblurring (Li et al., 2020; Dong et al., 2020; Zhang et al., 2022) and blind deconvolution (Krishnan et al., 2011). These methods are tailored to specific tasks. For example, Li et al. (2020) is recast from TV-regularization due to the sparse gradients in natural images, Dong et al. (2020) incorporates operations in Fourier space, which is hard to generalize to other inverse problem tasks. Their task-specific designs often involve numerous hyperparameters, leading to high sensitivity in reconstruction (Cho & Lee, 2009). For example, we observe the instability in training the blind deblurring algorithm in Li et al. (2020) to hyperparameters such as threshold value, parameter initialization, scale of proximity, etc. In contrast, we propose a simple but stable solution applicable to general IPs with only estimated forward models.
On the other hand, when initial and true forward models differ structurally, updating parameters in $\mathcal{A}_{0}$ often yields suboptimal solutions, common in many nonlinear inverse problems. Oftentimes, $\mathcal{A}_{0}$ is the best approximation within this structure. This motivates the use of a more complex structure, like a neural network, to approximate the true model. In nonlinear applications, reconstructions with erroneous model estimates are typically tackled in two steps. First, training a neural network to learn the true forward model (Candiani et al., 2021; Hauptmann et al., 2018; Koponen et al., 2021; Liu et al., 2013; Lunz et al., 2021). Then, solving for the reconstruction using a predetermined regularizer like TV-norm (Candiani et al., 2021) and pseudo-Huber function Lunz et al. (2021), or learning a regularizer using the pre-trained forward model (Koponen et al., 2021). These approaches assume either a “perfect” learned forward model Koponen et al. (2021), or errors can be corrected via the learned regularizer Hauptmann et al. (2018) which is less interpretable. In this work, we propose to use an untrained neural network to match the data consistency, which allows updating the forward model and the reconstruction in one step adaptively for each instance.
@@ 6 Experiments and Discussion
where $\bm{T}\in[0,1]^{3\times m\times n}$ is an unknown transmission map as a function of the fog-free image $\bm{x}$, $m$ and $n$ denote the width and height of the image. $\bm{T}(\bm{x})$ is then applied pointwise ($\odot$) to $\bm{x}$. $\bm{L}(\bm{x})$ is the unknown atmospheric light profile (Tufail et al., 2019). The goal is to estimate the model mismatch and recover the clean image. When the transition map is a matrix with all ones, the initial forward model is simplified to identity mapping pixel-wise. Since $\bm{T}$ and $\bm{L}$ are unknown functions with respect to $\bm{x}$, we use two separate networks to learn transmission mapping and the light profile respectively. This example shows the flexibility of our proposed methods. Instead of having one single network $f_{\theta}$ that learns the residual, if we know the explicit formulation of the true model, we can customize the residual network in a more structured way. To evaluate the proposed approaches, the CityScapes - Depth and Segmentation dataset given in Cordts et al. (2016) is used. This dataset provides depth maps, as well as foggy and clean views of urban street scenes.
@@ 6.1 Reconstruction Results
Table 1 compares the average and the standard deviation of the testing peak signal-to-noise ratio (PSNR) in dB and Structural Similarity Index Measure (SSIM) of the reconstructions. The neural network learns a direct inverse regardless of $\mathcal{A}_{0,i}$, which is the worst in reconstruction. It is notable that, even with some degree of inaccuracy in $\mathcal{A}$, the robust LU still outperforms the direct inverse mapping. This is because end-to-end training enables LU to fix the gradient-step errors resulting from the model mismatch to some extent, thereby maintaining reasonably higher-quality results. Our results align with the finding in Krainovic et al. (2023) that the robust LU trains an inverse mapping that is less sensitive to perturbations in the $\bm{y}$-space, by seeing inaccuracies in the forward model at train time. However, in the next section, we will discuss the intermediate reconstruction ineffectiveness of robust LU compared to the proposed methods.
@@
The proposed methods make an improvement by actively adjusting for the model mismatch for each data pair. Instead of having the neural network learn both the proximal step and the model correction as in a black box, the proposed methods separate the tasks and each part of the network is now learning a clearer and easier task, resulting in a substantial performance gain over both the robust LU and the direct inverse mapping in all tasks. Furthermore, $\mathcal{A}$-adaptive DEQ allows for a higher number of iterations than $\mathcal{A}$-adaptive LU, resulting in superior performance in both PSNR and SSIM. Notice that the $\mathcal{A}$-adaptive LU tends to have a smaller standard deviation than $\mathcal{A}$-adaptive DEQ because the latter varies in number of iterations.
Table 1: Average and standard deviation of testing PSNR and SSIM for direct inverse mapping, robust LU, $\mathcal{A}$-adaptive LU, and $\mathcal{A}$-adaptive DEQ. The best performances for each task are in bold and the second best performances are underlined.
@@
Direct Inverse PSNR 24.002 $\pm$ 0.101 22.816 $\pm$ 0.112 21.802 $\pm$ 1.706
SSIM 0.7689 $\pm$ 0.0007 0.7421 $\pm$ 0.0041 0.8738 $\pm$ 0.0347
Robust LU PSNR 34.380 $\pm$ 0.437 24.785 $\pm$ 0.836 29.645 $\pm$ 1.939
SSIM 0.9429 $\pm$ 0.0014 0.8407 $\pm$ 0.0214 0.9563 $\pm$ 0.0164
$\mathcal{A}$-adaptive LU PSNR 36.621 $\pm$ 0.330 27.271 $\pm$ 0.787 31.112 $\pm$ 1.854
(proposed meth. I) SSIM 0.9731 $\pm$ 0.0010 0.8960 $\pm$ 0.0157 0.9661 $\pm$ 0.0133
$\mathcal{A}$-adaptive DEQ PSNR 37.809 $\pm$ 0.532 28.236 $\pm$ 0.321 32.149 $\pm$ 2.069
(proposed meth. II) SSIM 0.9774 $\pm$ 0.0041 0.9082 $\pm$ 0.0201 0.9723 $\pm$ 0.0124
@@ 6.3 Intermediate Reconstructions of $\mathcal{A}$-adaptive DEQ
We show the intermediate reconstructions of the DEQ variant of the proposed method in Fig. 7. The proposed network is trained with a maximum of 30 iterations or when a fixed point solution is reached, whereas when evaluating $\mathcal{A}$-adaptive DEQ with more iterations (i.e., 100), the reconstruction error remains low. This result aligns with the finding in Gilton et al. (2021a) where the reconstruction quality remains high after reaching the fixed point, particularly when the forward model is precisely known.
@@ 6.4 Robustness of $\mathcal{A}$-adaptive LU with More Iterations
In this section, we empirically assess the stability of our proposed $\mathcal{A}$-adaptive LU variant by evaluating it with a greater number of unrolling iterations than those employed during training. The result can be found in 8. Across all three tasks, a noticeable Mean Squared Error (MSE) gap emerges between the two curves. The solid blue curve, representing robust LU, exhibits a drop in MSE at iteration 5, which corresponds to the target output during training. However, as the number of iterations exceeds the max iterations of training (in this set of experiments $k>5$), the MSE rises significantly. Notice that although the MSE at iteration 10 is relatively low for the landscape defogging task, the reconstruction is not stable and less predictable. In contrast, the dashed red curve ($\mathcal{A}$-adaptive LU) oscillates around the lowest MSE. This observation highlights the enhanced robustness of the proposed $\mathcal{A}$-adaptive LU when subjected to a higher number of iterations during evaluation.
@@ 7 Conclusion
This work adaptively addresses the errors in the forward model by optimizing the auxiliary variable and the untrained mismatch neural network within a model-based architecture. We show the flexibility of the proposed forward model matching module using two well-known architectures. However, the accuracy-runtime tradeoff still exists. Although the proposed $\mathcal{A}$-adaptive DEQ variant achieves the best performance for all tasks, when solving large-scale inverse problems, careful consideration of runtime is necessary. Algorithms such as momentum methods accelerate the convergence speed of an optimization problem. Investigating a more sophisticated design with accelerated algorithms offers an intriguing prospect for future work to expedite $\mathcal{A}$-adaptive DEQ.
Acknowledgments
This work was supported by the Center for Energy and Geo Processing (CeGP) at Georgia Tech, the Deanship of Research Oversight and Coordination (DROC), King Fahd University of Petroleum and Minerals (KFUPM), under Project GTEC2013.
Appendix A Appendix: Training Details
Appendix B Appendix: Effect of $\mathcal{A}_{0}$ in Reconstruction
In addition, we show that our proposed methods are less sensitive to the quality of the initial forward model. We demonstrate the idea using the image deblurring task, where one has flexibility in choosing the initial forward model. Assuming the true forward model for an image deblurring problem has a Gaussian kernel size of 5 with a variance of 7, we consider initial forward models $\bm{A}_{0}$ constructed with Gaussian kernels of sizes 5, but variances of 1, 3, 5, and 7. Figure 9 illustrates the reconstruction results with different $\bm{A}_{0}$’s for each method. When $\bm{A}_{0}$ deviates from $\bm{A}$, robust LU exhibits noticeable artifacts (with $\sigma=1$ and $3$), while our proposed methods show minimal degradation in quality even with significant differences between $\bm{A}_{0}$ and $\bm{A}$. This is due to the adaptive nature of our methods in iteratively addressing model mismatches. Notice that $\sigma=1$ and 3 are treated as extreme cases, which were never encountered in $\bm{A}$ and $\bm{A}_{0}$ during training. The proposed methods generalize better to unseen estimated forward models.
Figure 9: Visualizing the deblurred images using different $\bm{A}_{0}$, where $\sigma$ denotes the variance of the Gaussian kernel of the estimated forward model. The numbers in each image denote the PSNR and SSIM values.